OpenStudy (anonymous):

A toy store orders jigsaw puzzles for $3 and sells them for $4.50. What is the percent of markup? Round your answer to the nearest tenth of a percent, if necessary. A) 20% B) 33.3% C) 50% D) 66.7%

jimthompson5910 (jim_thompson5910):

First off, what is the profit?

OpenStudy (anonymous):

....

jimthompson5910 (jim_thompson5910):

Profit = (Amount Sold) - (Cost)

jimthompson5910 (jim_thompson5910):

For example, say you are a merchant and you buy a table for $100. Then you sell that table for $120. The profit would be... Profit = (Amount Sold) - (Cost) Profit = (120) - (100) Profit = 20 dollars

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

So what is the profit in this case?

jimthompson5910 (jim_thompson5910):

How much money does this merchant make from selling one puzzle?

OpenStudy (anonymous):

4.5-3=1.5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now divide that profit by the initial amount 3 1.5/3 = 0.5 Then multiply that result by 100 to convert it to a percentage 0.5*100 = 50% So the markup is 50%

jimthompson5910 (jim_thompson5910):

Notice how $1.50 is 50% of $3 So putting a 50% markup means that the seller earns back the full cost (of $3) and earns an extra 50% of the cost (which is $1.50)